OpenStudy (anonymous):

4m-4=4m solve for m show your work

OpenStudy (anonymous):

Solution for 4m-4=4m equation: Simplifying 4m + -4 = 4m Reorder the terms: -4 + 4m = 4m Add '-4m' to each side of the equation. -4 + 4m + -4m = 4m + -4m Combine like terms: 4m + -4m = 0 -4 + 0 = 4m + -4m -4 = 4m + -4m Combine like terms: 4m + -4m = 0 -4 = 0 Solving -4 = 0

OpenStudy (anonymous):

So Your answer would be -4=0! Need any more help?

OpenStudy (anonymous):

And here is another part of the answer this equation is invalid, the left and right sides are not equal, therefore there is no solution.

OpenStudy (neonumbrella5115):

4m-4=4m (Original equation) -4m -4m (Subtract 4m from each side) -4=0 Wait, but -4 does not equal 0. So that means that this equation has no solution

OpenStudy (anonymous):

@neonumbrella5115 Yea I could have made my equation smaller

OpenStudy (anonymous):

thanks, but solve for m

OpenStudy (anonymous):

That is M there is no Solution 4m-16<=4m-16 +16 +16 4m<=4m 4m/4 <= 4m/4 m <= m Since m will equal m for all values of m, there are infinite solutions. On a number line, this would just be a solid line extending into infinity in both directions

OpenStudy (anonymous):

So this is what it would look like on a graph/ Or number line 

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

No Problem! Need anymore help?

OpenStudy (anonymous):

Yeah can you do 12=-4(-6x-3)

OpenStudy (anonymous):

Yup hold on.

OpenStudy (anonymous):

Simplify? Right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Simplifying 12 = -4(6x + -3) x = 0

OpenStudy (anonymous):

Wait sorry i need all the work

OpenStudy (anonymous):

Ok Hold on

OpenStudy (anonymous):

Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '24x' to each side of the equation. 0 + 24x = -24x + 24x Remove the zero: 24x = -24x + 24x Combine like terms: -24x + 24x = 0 24x = 0 Divide each side by '24'. x = 0 Simplifying x = 0

OpenStudy (anonymous):

PM me for anymore questions.

OpenStudy (neonumbrella5115):

@BraydonG It's all good.